OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±6) and foci at (0, ±9).

OpenStudy (amistre64):

do you recall the general equation for a hyperbola ?

OpenStudy (anonymous):

\[\frac{(x-h)^{2} }{ a ^{2} }-\frac{(y-k)^{2} }{ b ^{2} }\]

OpenStudy (amistre64):

... yeah, and =1 , can you tell me from the given information where the center of this will be?

OpenStudy (anonymous):

its at origin, so itll be, x^2/a^2 - y^2/b^2 =1

OpenStudy (anonymous):

sorry its easy to type than use the equation tool

OpenStudy (amistre64):

good, and one more bit of information from the givens .... will x or y be after the subtraction sign?

OpenStudy (anonymous):

if its in standard for, x should be before the subtraction sign

OpenStudy (amistre64):

standard for a hyperbola can be defined 2 different ways due to the subtraction operation: we are given that vertexes that when x=0, y is some value: (0/a)^2 - (y/b)^2 = 1 - (y/b)^2 = 1 is not possible among the real numbers, so we need to subtract the x parts instead of the y parts (y/b)^2 - (x/a)^2 = 1; now this is doable

OpenStudy (amistre64):

using the vertex again; we are able to define the b value .... can you tell me what b has to be?

OpenStudy (anonymous):

um, well i know that c^2 = a^2 + b^2 so i can use that to find b b^2=c^2=a^2 and c is the distance that the foci are from the center, which in this case is origin

OpenStudy (anonymous):

sorry one sec

OpenStudy (anonymous):

b^2=c^2-a^2

OpenStudy (amistre64):

that was going to be my next step :) but we can determine the value for b before we go thru that; and will use it in the process to find a

OpenStudy (anonymous):

so c=9 b^2= 81-a^2

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the vertex tells us the when x=0, y = +-6 ( |6|/b)^2 - (0/a)^2 = 1 (6/b)^2 = 1 ; what is b now?

OpenStudy (anonymous):

(6/b)^2 = 1 6/b = sqrt(1) 6/b = 1 6 = b

OpenStudy (amistre64):

good, now we can use that to determine "a" :)

OpenStudy (anonymous):

okay so, 81 = a^2 + 36 a^2 = 81-36 a^2 = 45 a= sqrt(45)

OpenStudy (anonymous):

3sqrt(5)

OpenStudy (amistre64):

we can usually stop at a^2 = 45 since its asking for the standard equation and not some distances: this gives us the completion as:\[\frac{y^2}{36}-\frac{x^2}{45}=1\]

OpenStudy (anonymous):

thank you for all the help :D

OpenStudy (amistre64):

youre welcome :)